The demand curve for the bottle of water is Qd=100-6P and the supply Qs=28+3P.what is the equilibrium price?

1
Expert's answer
2022-05-02T13:40:37-0400

"Q_d=100-6P,"

"Q_s=28+3P,"

"Q_d=Q_s,"

"100-6P=28+3P,"

"72=9P,"

"P=8,"

"Q_d=Q_s=52."